OpenStudy (anonymous):

The perimeter of a rectangle is 52. The length is 6 greater than the width. Find the length and width of this rectangle.

OpenStudy (anonymous):

someeone answer it

OpenStudy (anonymous):

answer it Mybrainvsme

OpenStudy (anonymous):

working on it.

OpenStudy (anonymous):

I wish i had paper

OpenStudy (anonymous):

ok got it hold on

OpenStudy (anonymous):

:D w = 10 and L = 16. 10 + 10 +16 +16 = 52

OpenStudy (anonymous):

L = w+ 6 so plug this into 2w+2l = 52 2w+2(w+6) = 52 2w+2w+12 = 52 2w+2w = 40 4w=40 w = 10 l = 10 + 6 l = 16

OpenStudy (anonymous):

sure lol 2(x+6)+2y=52 simplify

OpenStudy (anonymous):

math don't you take my glory.

OpenStudy (anonymous):

i'm fluttering hungry.